OpenStudy (anonymous):

Find the measures of angle R and angle N. Given:Triangle RST is congruent to triangle NPQ, the measure of angle R = -7x + 9, and the measure of angle N = -10x

Directrix (directrix):

Given:Triangle RST is congruent to triangle NPQ. This means that <R ≅ <N because correspoinding angles of congruent triangles are congruent. So, -7x + 9 = -10x 9 = -3x -3 = x m<R = -7*(-3) + 9 = ? m<N = -10*(-3) = ? @Anonymous120 Note that the work and answers posted by @calculusxy are incorrect. In Euclidean Geometry, angle measures are never negative.